# Relative equilibria for Scaling Symmetries and Central Configurations

### v1

@@ 5.2 Properties of Conformal Momentum Maps
Let $\Psi$ be an action of $G=\mathbb{R}^{+}$ on $Q$, and let $\Phi=\hat{\Psi}^{c}$ be the scaled cotangent lifted action with parameter $c$. Then the conformal momentum map of $\hat{\Psi}^{c}$ is conformally invariant.
@@
Given the action $\Phi=\hat{\Psi}^{c}$, consider the associate diffeomorphism $\Phi_{g}=\hat{\Psi}_{g}^{c}:T^{\ast}Q\to T^{\ast}Q$. Then for any $g\in G$ we have
@@
$=\left\langle\hat{\Psi}_{g}^{c}(\alpha_{q}),\xi_{Q}(\Psi_{g}(q))\right\rangle$
$=\left\langle\alpha_{q},g^{c}\,T\Psi_{g^{-1}}\xi_{Q}(\Psi_{g}(q))\right\rangle\quad\mbox{By definition of }\hat{\Psi}_{g}^{c}$
$=g^{c}\left\langle\alpha_{q},T_{\Psi_{g}(q)}\Psi_{g^{-1}}\circ\xi_{Q}\circ\Psi_{g}(q))\right\rangle$
$=g^{c}\left\langle\alpha_{q},(\Psi^{\ast}_{g}\xi_{Q})(q)\right\rangle\quad\mbox{by definition of pullback}$
@@
$=\left\langle g^{c}\operatorname{Ad}^{\ast}_{g^{-1}}J(\alpha_{q}),\xi\right\rangle$
$=\left\langle g^{c}J(\alpha_{q}),\xi\right\rangle.$
Hence $J(\Psi_{g}(\alpha_{q}))=g^{c}J(\alpha_{q})$. ∎
@@ 6 Relative Equilibria of Scaling Symmetries
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold. Suppose there is a conformal symplectic action $\Phi$ of the multiplicative group $G=\mathbb{R}^{+}$ on $M$. Assume that $H$ is conformally invariant under $\Phi$, implying that $\Phi$ represents a scaling symmetry of the Hamiltonian system.
@@
Then $z_{e}\in M$ is called a relative equilibrium (with respect to the scaling symmetry) of the Hamiltonian vector field $X_{H}$ if, for every $t\in I$, there exists $\eta(t)\in G$ such that $\eta(0)=\operatorname{Id}$ and $\gamma(t)=\Phi_{\eta(t)}(z_{e})$.
@@
    The curve $\gamma(t)$, the maximal integral curve of the Hamiltonian system with $\gamma(0)=z_{e}$, is contained in the group orbit $\operatorname{Orb}(z_{e})=\{\Phi_{g}(z_{e})\,|\,g\in G\}$.
@@
    There exists an element $\xi$ in the Lie algebra $\mathfrak{g}$ of the symmetry group $G$ such that:
@@
    There exists $\xi\in\mathfrak{g}$ such that $\gamma(t)=\Phi_{\eta(t)}(z_{e})$ is a maximal integral curve of the Hamiltonian system $(M,\omega,H)$ with
@@
$(i)\implies(ii)$. By definition, for each $t\in I$ there exists $\eta(t)\in G$ such that $\eta(0)=\operatorname{Id}$ and $\gamma(t)=\eta(t)\cdot z_{e}=\Phi_{\eta(t)}(z_{e})$. Thus, $\gamma(t)$ remains in the group orbit of $z_{e}$ for all $t\in\mathbb{R}$.
$(ii)\implies(i)$. Let $\gamma:I\to\mathbb{R}$ be the maximal integral curve of the Hamiltonian system with $\gamma(0)=z_{e}$. Then, by assumption, $\gamma(t)$ lies in $\operatorname{Orb}(z_{e})$, for all $t\in I$. Hence, by definition of the group orbit, for each $t\in\mathbb{R}$ there is $\eta(t)\in G$ such that
@@
$(i)\implies(iii)$. By the definition of relative equilibrium it follows that the trajectory $\gamma(t)$ through $z_{e}$ is contained in a single group orbit $\operatorname{Orb}(z_{e})$, so its derivative $X_{H}(z_{e})=\dot{\gamma}(0)$ lies on the tangent space to the group orbit $T_{z_{e}}(\operatorname{Orb}(z_{e}))=\{\xi_{M}(z_{e})|\xi\in\mathfrak{g}\}$, where $\xi_{M}(z_{e})$ is the infinitesimal generator of the action at $z_{e}$. Hence, there is $\xi\in\mathfrak{g}$ such that $X_{H}(z_{e})=\xi_{M}(z_{e})$.
@@
First, let $y(t)=\Phi(\exp(t\xi),z_{e})$, where $\Phi:G\times M\to M$ is the group action. Clearly, $y(0)=z_{e}$. We want to show that $y(t)$ satisfies the ODE $\frac{dy}{dt}=\xi_{M}(y(t))$. Since $\xi_{M}(z_{e})=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi(\exp(\tau\xi),z_{e})$, we have that if $s=t+\tau$ then
@@
so $y(t)$ is a path satisfying $y(0)=z_{e}$ and satisfying the ODE $\frac{dy}{dt}=\xi_{M}(y(t))$.
Next we want to show that $\xi_{M}(y(t))=T\Phi_{\exp(t\xi)}(\xi_{M}(z_{e}))$. Recall that if $f:M\to N$ is a map then the tangent map $(T_{x})f(v)=\left.\frac{d}{d\tau}\right|_{\tau=0}f(\alpha(\tau))$, where $\alpha(\tau)$ is a path such that $\alpha(0)=x$ and $\left.\frac{d}{d\tau}\right|_{\tau=0}\alpha(\tau)=v$, where $T_{x}f=f_{\ast}(x)$. From this definition it is easy to see that
@@
$=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{\exp(t\xi)}(y(\tau))=T\Phi_{\exp(t\xi)}(\xi_{M}(z_{e}))$
We now seek a formula for $X_{H}(y(t))$. Since the action is conformally symplectic with parameter $c$, and $H$ is conformally invariant with parameter $b$, we have by Proposition 4.5 that for each $g\in G$
$X_{H}(\Phi_{g}(z_{e}))=(g^{b-c})\cdot(\Phi_{g})_{\ast}(X_{H}(z_{e}))=(g^{b-c})\cdot T\Phi_{g}(X_{H}(z_{e})).$
Then, taking $g=\exp(\xi t)$ we obtain
@@
$=\exp((b-c)t)\cdot T\Phi_{\exp\xi t}(X_{H}(z_{e}))$
$=\exp((b-c)t)\cdot T\Phi_{\exp\xi t}(\xi_{M}(z_{e}))$
$=\exp((b-c)t)\cdot\xi_{M}(y(t))$
@@
If $b=c$, then using equation (10) and (12), we see that
@@
Hence, the curves $y(t)$ and $\gamma(t)$ are maximal integral curves of the same differential equations, which satisfy the same initial condition. It follows that $y(t)$ and $\gamma(t)$ coincide, so that
$\gamma(t)=\Phi_{\exp(t\xi)}(z_{e})=\Phi_{\eta(t)}(z_{e})=e^{\xi t}z_{e},$
and $\eta(t)=e^{\xi t}$.
If $b\neq c$, on the other hand, using equation (10) and (12) we obtain
$\frac{d}{dt}y(t)=\xi_{M}(y(t))=\exp\left((c-b)\xi t\right)\cdot X_{H}(y(t)).$
Letting $\gamma(\tau)=y(t(\tau))$ and performing the change of time $\frac{d\tau}{dt}=\exp((c-b)\xi t)$ yields
$\frac{d}{d\tau}\gamma(\tau)=X_{H}(\gamma(\tau)).$
Solving $\frac{d\tau}{dt}=\exp((c-b)\xi t)$ with the condition $\tau(0)=0$ yields
$\tau=\frac{1}{(c-b)\xi}\left[-1+\exp(\xi(c-b)t)\right].$
Solving for $t$ we obtain
$t=\frac{1}{\xi}\ln\left[(c-b)\xi\tau+1\right]^{1/(c-b)}$
and hence
$\gamma(\tau)=\Phi_{\exp(\xi t(\tau))}(z_{e})=\Phi_{\eta(\tau)}(z_{e})$
with $\eta(\tau)=[(c-b)\xi\tau+1]^{1/(c-b)}$.
@@
Hence $X_{H}(z_{e})=\xi_{M}(z_{e})$ for some $\xi\in\mathfrak{g}$ such that $\xi=\frac{d}{dt}g(0)$.
@@
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold. Suppose there is a conformal symplectic action $\Phi$ of the multiplicative group $G=\mathbb{R}^{+}$ on $M$. Assume that $H$ is conformally invariant under $\Phi$, and that $J$ is the conformal momentum map with parameter $c$ of this action. Then the point $z_{e}$ is a relative equilibrium if and only if there is a $\xi\in\mathfrak{g}$ such that
@@
If, on the other hand, $dH_{\xi}+(c\xi)\theta=0$, then by reversing the computations above and using the nondegeneracy of the symplectic form $\omega$, it follows that $X_{H}-X_{J_{\xi}}^{\xi c}=0$. This implies, by the definition of the conformal momentum map, that $X_{H}$ points in the direction of the group orbit through $z_{e}$.
@@
Let $(T^{\ast}Q,\omega_{0},\theta_{0},H)$ be a Hamiltonian system on $T^{\ast}Q$. Let $\Psi$ be an action of the multiplicative group $G=\mathbb{R}^{+}$ on $Q$, and let $\Phi=\hat{\Psi}^{c}$ be the induced scaled cotangent lifted action on $T^{\ast}Q$ (this action is conformally symplectic). Assume that $H$ is conformally invariant under $\Phi$. It follows that $\Phi$ is a scaling symmetry.
Let $z_{e}=(q_{e},p_{e})$ denote a covector based at $q_{e}\in Q$. If $z_{e}=(q_{e},p_{e})$ is a relative equilibrium for the scaling symmetry, then $q_{e}$ is called a central configuration.
@@
In studying this problem, we assume that the center of mass of the particles is at the origin, i.e., $m_{1}q_{1}+m_{2}q_{2}+\cdots+m_{n}q_{n}=0$. This assumption simplifies the analysis by avoiding technical complications arising from the non-trivial action of the scaling group $G=\mathbb{R}^{+}$ on the translation group $\mathbb{R}^{2}$. In fact, translations and scalings together form a semidirect product.
Consider the action of the group $G=\mathbb{R}^{+}$ on $T^{\ast}Q$ given by
@@
From Example 5.5, we know that $\Phi$ is a scaled cotangent lift, implying that $\Phi$ is a conformally symplectic action. Furthermore, it is easy to verify that $\Phi^{\ast}H=g^{-1}H$, showing that $\Phi$ is a scaling symmetry, sometimes referred to as Kepler scaling. The momentum map for this action, as noted in Example 5.5, is $J=\mathbf{p}^{T}\mathbf{q}$.
@@
$\frac{\partial H}{\partial\mathbf{q}}d\mathbf{q}$ $+\frac{\partial H}{\partial\mathbf{p}}d\mathbf{p}-\xi\left(\frac{\partial J}{\partial\mathbf{q}}d\mathbf{q}+\frac{\partial J}{\partial\mathbf{p}}d\mathbf{p}\right)+c\xi\mathbf{p}d\mathbf{q}$

### v2

@@ 5.2 Properties of Conformal Momentum Maps
Let $\Psi$ be an action of the multiplicative group $\mathbb{R}^{+}$ on $Q$, and let $\hat{\Psi}^{c}$ be the scaled cotangent lift of this action with parameter $c$. Then, the conformal momentum map associated with $\hat{\Psi}^{c}$ is conformally invariant.
@@
Consider the diffeomorphism $\hat{\Psi}_{g}^{c}:T^{*}Q\to T^{*}Q$ associated with the scaled cotangent lift action $\hat{\Psi}^{c}$. Let $z=(q,p)\in T^{*}Q$, where $q=\pi(z)$ is the base point under the canonical projection $\pi:T^{*}Q\to Q$, and $p\in T^{*}_{q}Q$ is the corresponding covector. Then, since
$\hat{\Psi}_{g}^{c}(z)=(\Psi_{g}(q),g^{c}\cdot(d\Psi_{g^{-1}})^{*}(p)),$
for any $g\in\mathbb{R}^{+}$, we have the following:
@@
$=\left\langle p,g^{c}\,(d\Psi_{g^{-1}})\xi_{Q}(\Psi_{g}(q))\right\rangle$
$=g^{c}\left\langle p,d\Psi_{g^{-1}}\circ\xi_{Q}\circ\Psi_{g}(q)\right\rangle$
@@
[…]
$=\left\langle g^{c}\operatorname{Ad}^{*}_{g^{-1}}J(z),\xi\right\rangle$
$=\left\langle g^{c}J(z),\xi\right\rangle.$
Thus, we conclude that $J(\hat{\Psi}_{g}(z))=g^{c}J(z)$, confirming the conformal invariance of $J$. ∎
@@ 6 Relative Equilibria of Scaling Symmetries
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold. Suppose there is a conformal symplectic action $\Phi$ of the multiplicative group $\mathbb{R}^{+}$ on $M$. Assume that $H$ is conformally invariant under $\Phi$, implying that $\Phi$ represents a scaling symmetry of the Hamiltonian system.
@@
Then $z_{e}\in M$ is called a relative equilibrium (with respect to the scaling symmetry) of the Hamiltonian vector field $X_{H}$ if, for every $t\in I$, there exists $\eta(t)\in\mathbb{R}^{+}$ such that $\eta(0)=\operatorname{Id}$ and $\gamma(t)=\Phi_{\eta(t)}(z_{e})$.
@@
    The curve $\gamma(t)$, the maximal integral curve of the Hamiltonian system with $\gamma(0)=z_{e}$, is contained in the group orbit $\operatorname{Orb}(z_{e})=\{\Phi_{g}(z_{e})\,|\,g\in\mathbb{R}^{+}\}$.
@@
    There exists an element $\xi$ in the Lie algebra $\mathbb{R}$ of the symmetry group $\mathbb{R}^{+}$ such that:
@@
    There exists a Lie algebra element $\xi\in\mathbb{R}$ such that $\gamma(t)=\Phi_{\eta(t)}(z_{e})$ is a maximal integral curve of the Hamiltonian system $(M,\omega,H)$ with
@@
The logic of the proof will go as follows:
$(i)\iff(ii),\quad\mbox{and}\quad(i)\implies(iii)\implies(iv)\implies(i).$
$(i)\implies(ii)$. By definition, for each $t\in I$ there exists $\eta(t)\in\mathbb{R}^{+}$ such that $\eta(0)=\operatorname{Id}$ and $\gamma(t)=\Phi_{\eta(t)}(z_{e})$. Thus, $\gamma(t)$ remains in the group orbit of $z_{e}$ for all $t\in\mathbb{R}$.
$(ii)\implies(i)$. Let $\gamma:I\to\mathbb{R}$ be the maximal integral curve of the Hamiltonian system with $\gamma(0)=z_{e}$. Then, by assumption, $\gamma(t)$ lies in $\operatorname{Orb}(z_{e})$, for all $t\in I$. Hence, by definition of the group orbit, for each $t\in\mathbb{R}$ there is $\eta(t)\in\mathbb{R}^{+}$ such that
@@
$(i)\implies(iii)$. By the definition of relative equilibrium it follows that the trajectory $\gamma(t)$ through $z_{e}$ is contained in a single group orbit $\operatorname{Orb}(z_{e})$, so its derivative $X_{H}(z_{e})=\dot{\gamma}(0)$ lies on the tangent space to the group orbit $T_{z_{e}}(\operatorname{Orb}(z_{e}))=\{\xi_{M}(z_{e})|\xi\in\mathbb{R}\}$, where $\xi_{M}(z_{e})$ is the infinitesimal generator of the action at $z_{e}$. Hence, there is $\xi\in\mathbb{R}$ such that $X_{H}(z_{e})=\xi_{M}(z_{e})$.
@@
First, consider the path $y(t)=\Phi(\exp(t\xi),z_{e})$, where $\Phi:\mathbb{R}^{+}\times M\to M$ is the group action. Since $y(0)=z_{e}$, we want to verify that $y(t)$ satisfies the ODE $\frac{dy}{dt}=\xi_{M}(y(t))$. The infinitesimal generator $\xi_{M}(z_{e})$ is given by:
$\xi_{M}(z_{e})=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi(\exp(\tau\xi),z_{e}).$
Then, using a shift in the time parameter $s=t+\tau$, we have:
@@
so $y(t)$ is a path satisfying $y(0)=z_{e}$ and the ODE $\frac{dy}{dt}=\xi_{M}(y(t))$.
Next, we want to show that $\xi_{M}(y(t))=d\Phi_{\exp(t\xi)}(\xi_{M}(z_{e}))$. Recall that for a smooth map $f:M\to N$, the differential $df$ acts on a vector $v\in T_{x}M$ as:
$df_{x}(v)=\left.\frac{d}{d\tau}\right|_{\tau=0}f(\alpha(\tau)),$
where $\alpha(\tau)$ is a path in $M$ such that $\alpha(0)=x$ and $\left.\frac{d}{d\tau}\right|_{\tau=0}\alpha(\tau)=v$. The differential $df_{x}$ is also often denoted by $f_{\ast}(x)$ and called pushforward. Using this definition, we obtain:
@@
$=\left.\frac{d}{d\tau}\right|_{\tau=0}\Phi_{\exp(t\xi)}(y(\tau))$
$=d\Phi_{\exp(t\xi)}(\xi_{M}(z_{e})).$
We now seek a formula for $X_{H}(y(t))$.
Since the action is conformally symplectic with parameter $c$ and $H$ is conformally invariant with parameter $b$, it follows by Proposition 4.5 that
$((\Phi_{g})^{\ast}X_{H})(z_{e})=g^{b-c}X_{H}(\Phi_{g}(z_{e})).$
Applying $(\Phi_{g})_{\ast}$ to both sides of this equation, we obtain, for each $g\in\mathbb{R}^{+}$,
$(\Phi_{g})_{\ast}\left(((\Phi_{g})^{\ast}X_{H})(z_{e})\right)=(\Phi_{g})_{\ast}\left(g^{b-c}X_{H}(\Phi_{g}(z_{e}))\right).$
Since $(\Phi_{g})_{\ast}\circ(\Phi_{g})^{\ast}$ acts as the identity, this simplifies to
$X_{H}(\Phi_{g}(z_{e}))=g^{b-c}\cdot((\Phi_{g})_{\ast}X_{H})(\Phi_{g}(z_{e}))=g^{b-c}\cdot d\Phi_{g}(X_{H}(z_{e})),$
where we have expressed the pushforward as $d\Phi_{g}$ for clarity.
Then, taking $g=\exp(\xi t)$ we find
@@
$=\exp((b-c)\xi t)\cdot d\Phi_{\exp\xi t}(X_{H}(z_{e}))$
$=\exp((b-c)\xi t)\cdot d\Phi_{\exp\xi t}(\xi_{M}(z_{e}))$
$=\exp((b-c)\xi t)\cdot\xi_{M}(y(t))$
@@
If $b=c$, then using equations (10) and (12), we find that
@@
Since $\gamma(t)$ is an integral curve of the Hamiltonian system, we must have $\frac{d}{dt}\gamma(t)=X_{H}(\gamma(t))$. Therefore, the curves $y(t)$ and $\gamma(t)$ are maximal integral curves of the same differential equation, with the same initial condition. Consequently, $y(t)$ and $\gamma(t)$ coincide, which implies that
$\Phi_{\exp(t\xi)}(z_{e})=y(t)=\gamma(t)=\Phi_{\eta(t)}(z_{e}),$
and thus $\eta(t)=e^{\xi t}$.
If $b\neq c$, we need to show that $\gamma(t)$, the Hamiltonian integral curve, is a reparametrization of the group trajectory $y(t)$. Using equation (10) and (12) we obtain
$\frac{d}{dt}y(t)=\xi_{M}(y(t))=\exp\left((c-b)\xi t\right)\cdot X_{H}(y(t)),$
while $\gamma(t)$ to be an integral curve of the Hamiltonian system we must have $\frac{d}{d\tau}\gamma(\tau)=X_{H}(\gamma(\tau))$. This shows that the group trajectory $y(t)$ is not directly an integral curve of the Hamiltonian vector field $X_{H}$, but rather follows a scaled version of it, with the scaling factor $\exp((c-b)\xi t)$.
To convert $y(t)$ into an actual Hamiltonian trajectory, we introduce a new time variable $\tau$, so that $\gamma(\tau)=y(t(\tau))$ satisfies the equation:
$\frac{d}{d\tau}y(t(\tau))=X_{H}(y(t(\tau))).$
We can achieve this by defining $\tau$ such that:
$\frac{d\tau}{dt}=\exp((c-b)\xi t).$
Solving this differential equation with the initial condition $\tau(0)=0$ gives:
$\tau=\frac{1}{(c-b)\xi}\left[\exp\left((c-b)\xi t\right)-1\right].$
Now, inverting this expression to find $t$ in terms of $\tau$, we get:
$t=\frac{1}{(c-b)\xi}\ln\left[(c-b)\xi\tau+1\right].$
Thus, we can express the integral curve $\gamma(\tau)$ as a reparametrization of $y(t)$ using $t(\tau)$:
$\gamma(\tau)=y(t(\tau))=\Phi_{\exp(\xi t(\tau))}(z_{e}).$
Substituting $t(\tau)$ into the expression for $\gamma(\tau)$ given above, we obtain:
$\gamma(\tau)=\Phi_{\eta(\tau)}(z_{e}),\quad\text{where}\quad\eta(\tau)=\left[(c-b)\xi\tau+1\right]^{1/(c-b)}.$
This shows that $\gamma(\tau)$ is indeed an integral curve of the Hamiltonian system, reparametrized from the scaled group trajectory $y(t)$.
$(iv)\implies(i)$. This is clear, since we are given an explicit formula for $\gamma(t)=\Phi_{\eta(t)}(z_{e})$.
@@
Hence $X_{H}(z_{e})=\xi_{M}(z_{e})$ for some Lie algebra element $\xi\in\mathbb{R}$ such that $\xi=\frac{d}{dt}g(0)$.
@@
Let $(M,\omega,\theta,H)$ be a Hamiltonian system on an exact symplectic manifold. Suppose there is a conformal symplectic action $\Phi$ of the multiplicative group $\mathbb{R}^{+}$ on $M$. Assume that $H$ is conformally invariant under $\Phi$, and that $J$ is the conformal momentum map with parameter $c$ of this action. Then the point $z_{e}$ is a relative equilibrium if and only if there is a Lie algebra element $\xi\in\mathbb{R}$ such that
@@
If, on the other hand, $dH_{\xi}+(c\xi)\theta=0$, then by reversing the computations above and using the nondegeneracy of the symplectic form $\omega$, it follows that $X_{H}-X_{J_{\xi}}^{\xi c}=0$. This implies, by the definition of the conformal momentum map, that $X_{H}=\xi_{M}$, that is, $X_{H}$ points in the direction of the group orbit through $z_{e}$.
@@
Let $(T^{\ast}Q,\omega_{0},\theta_{0},H)$ be a Hamiltonian system on $T^{\ast}Q$. Let $\Psi$ be an action of the multiplicative group $\mathbb{R}^{+}$ on $Q$, and let $\Phi=\hat{\Psi}^{c}$ be the induced scaled cotangent lifted action on $T^{\ast}Q$, which is conformally symplectic. Assume that $H$ is conformally invariant under $\Phi$ so that $\Phi$ is a scaling symmetry of the system.
Now, let $z_{e}=(q_{e},p_{e})$ denote a covector based at $q_{e}\in Q$. If $z_{e}=(q_{e},p_{e})$ is a relative equilibrium for the scaling symmetry, then the point $q_{e}\in Q$ is called a central configuration.
@@
We assume that the center of mass of the particles is located at the origin, i.e., $m_{1}q_{1}+m_{2}q_{2}+\cdots+m_{n}q_{n}=0$. This assumption simplifies the analysis by eliminating technical complications due to the non-trivial interaction between the scaling group $\mathbb{R}^{+}$ and translations in $\mathbb{R}^{2}$. Together, translations and scalings form a semidirect product, so fixing the center of mass at the origin allows us to isolate the effects of scaling.
Consider the action of the group $\mathbb{R}^{+}$ on $T^{\ast}Q$ given by
@@
From Example 5.5, we know that $\Phi$ is a scaled cotangent lift, implying that $\Phi$ is a conformally symplectic action. Furthermore, it is easy to verify that $\Phi^{\ast}H=g^{-1}H$, showing that $\Phi$ is a scaling symmetry, sometimes referred to as Kepler scaling [3]. The momentum map for this action, as noted in Example 5.5, is $J=\mathbf{p}^{T}\mathbf{q}$.
@@
$d(H-\xi J)+(c\xi)\theta_{0}$ $=\frac{\partial H}{\partial\mathbf{q}}d\mathbf{q}+\frac{\partial H}{\partial\mathbf{p}}d\mathbf{p}-\xi\left(\frac{\partial J}{\partial\mathbf{q}}d\mathbf{q}+\frac{\partial J}{\partial\mathbf{p}}d\mathbf{p}\right)+c\xi\mathbf{p}d\mathbf{q}$